Three objects, O2, O1 and O3 are charged and O2, O1 have equal charges. Suppose the flux through the surface A1that encloses all three objects is 4.40 Three objects, O2, O1 and O3 are charged 105 N
Answer #1

apply electric flux phi = EA cos theta   = Q/eo


ae flux 3 = q3/eo

q3 = O3 = 2.6 e5 * 8.85 e-12

q3   = O3 = 2.301 uC

---------------------

q1= q2 = q

so Qnet = 2q

so ( 2q + 2.301uC) = 4.4 e 5 * 8.85 e-12

2q = 3.89 uC

2q = 1.589 uC

q1 = q2 = 0.795 uC

q3 = 2.301 uC